OpenStudy (anonymous):

a circle has a radius of 6.7 cm. a) find the circumference of the circle . round your answer to the nearest centimeter. b) find the area of the circle. round your to the nearest square centimeter

OpenStudy (anonymous):

circumfrance of a circle = 2 * pi * radius we have a radius of 6.7, so the equation becomes = 2 * pi * 6.7 can you solve for the area now?

OpenStudy (anonymous):

circumfrance is 42.076 @DemolisionWolf

OpenStudy (anonymous):

good number, yep, not just round it so there arn't any decimal numbers

OpenStudy (anonymous):

how do i round it to the nearest whole centimeter? @DemolisionWolf and how do i solve for the area?

OpenStudy (anonymous):

rounding to the whole centimeter means no decimal numbers so, to round, we look at the numbers in the decimal area, if it is 5 or greater, we make the number to it's left +1 lets begin: 42.076 6 is > than 5, so 7+1 = 8 42.08 8 is > 5, so 0+1 = 1 42.1 1 is < 5, so nothing 42 42 is the anwer

OpenStudy (anonymous):

area = pi * r^2

OpenStudy (anonymous):

r means radius right? @DemolisionWolf